# RoundTable: Investigating Group Decision-Making Mechanism in Multi-Agent Collaboration

### v1

@@ Abstract
This study investigates the efficacy of Multi-Agent Systems in eliciting cross-agent communication and enhancing collective intelligence through group decision-making in a decentralized setting. Unlike centralized mechanisms, where a fixed hierarchy governs social choice, decentralized group decision-making allows agents to engage in joint deliberation. Our research focuses on the dynamics of communication and decision-making within various social choice methods. By applying different voting rules in various environments, we find that moderate decision flexibility yields better outcomes. Additionally, exploring the linguistic features of agent-to-agent conversations reveals indicators of effective collaboration, offering insights into communication patterns that facilitate or hinder collaboration. Finally, we propose various methods for determining the optimal stopping point in multi-agent collaborations based on linguistic cues. Our findings contribute to a deeper understanding of how decentralized decision-making and group conversation shape multi-agent collaboration, with implications for the design of more effective MAS environments.
@@ 1 Introduction
Collaboration is a fundamental aspect of the nature and human society. Whether among humans or animals, working together allows groups to overcome individual limitations and achieve greater collective outcomes. In nature, collaboration often arises as a strategy to boost survival, enhance resource gathering, or increase efficiency in completing tasks (Schmidt & Mech, 1997). Similarly, in human societies, collaboration drives innovation, facilitates problem-solving, and fosters shared understanding, enabling individuals to address complex challenges that would be unmanageable otherwise (De Man & Duysters, 2005; Graesser et al., 2018; Bittner & Leimeister, 2013). This innate tendency to collaborate is evident across various domains, from social communities to technological systems, where multiple entities coordinate their efforts toward a common goal. As we advance in developing intelligent agents, understanding and replicating these collaborative dynamics in artificial systems has become increasingly important, predominantly to cope with the complexity and adaptability seen in real-world interactions.
Agents powered by Large Language Models (LLMs) have demonstrated impressive problem-solving capabilities across a wide range of tasks. However, single-agent systems encounter significant difficulties when tasked with problems that are either too large or complex, often resulting in instability, misalignment with the intended request, and hallucination (Liu et al., 2024; Kuhn et al., 2023; Lyu et al., 2023). To address these limitations, research has increasingly turned toward Multi-Agent Systems (MAS). MAS have shown greater efficacy in harnessing collective intelligence by allowing individual agents to specialize in distinct skills and facilitating effective collaboration among them (Guo et al., 2024).
When agents working together in a MAS, it is natural for them to have varying interpretations and perspectives. While some opinions may align, disagreements are also frequent. This creates an inevitable tension between cooperation and competition, stemming from differences in backgrounds, information access, and individual goals. Therefore, the process of aggregating models’ diverse predictions into a final group decision becomes a crucial aspect of dynamic multi-agent collaboration.
Many existing LLM-driven MAS are designed based on centralized group decision-making, which typically involves layered or centralized architectures with a hierarchy. When there is a conflict between agents, it is resolved by a pre-assigned agent or a pre-defined process. These systems are often used for tasks where agent networks follow the waterfall method, leading to a stratified arrangement among agents (Hong et al., 2023; Qian et al., 2023; Dong et al., 2023). However, this hierarchical setup poses several critical challenges: (1) fairness: individual agent messages may not be accurately represented in the final group outcome, leading to potential misrepresentation (Jiang & Lu, 2019); (2) rigidity: the system’s fixed structure may lead to over-fitting to specific scenarios (Chen et al., 2023), reducing its adaptability to diverse and dynamic environments; and (3) bias: the agent with final decision-making power may introduce its own biases into the process, potentially distorting the outcomes (Owens et al., 2024). Additionally, centralized MAS cannot perform in environments requiring independent agent decisions, where private or incomplete information hinders a central authority from dictating optimal strategies (Xu et al., 2023).
@@
Figure 1: Overview of our multi-agent collaboration platform: RoundTable. It uses a round-based collaboration where agents simultaneously send messages, propose solutions, and vote. Based on a social choice mechanism, RoundTable selects the most preferred proposal for group decisions. Brief introduction of RoundTable is in Section 3.1, details are in Appendix A.1.
Decentralized group decision-making can ease these issues by distributing power among agents, where each agent has the ability to participate in the process. This is a common setting in world simulation and embodied environment, where agents need to behave independently because there exists information asymmetry or data boundary between agents (Mandi et al., 2024; Zhang et al., 2023a; Xu et al., 2023; Park et al., 2023), and possibly variability in capabilities that different agents have. With decisions made by multiple agents, the flexible structure of decentralized MAS adapts to various environments, but the lack of a fixed hierarchy demands careful monitoring of collaboration patterns.
Centralized and decentralized group decision-making attempt to mimic the ways in which human societies form collective policies, such as in monarchies and democracies. These decision-making mechanisms, known as social choice, are studied across various fields including economics, mathematics, philosophy, and social science, with the goal of aggregating and synthesizing individual preferences into a unified consensus. However, the impact of social choice methods on LLM-based MAS has yet to be explored.
In this study, we evaluate various social choice methods across different environments to observe and analyze agents’ group behaviors and collaboration pattern. This paper aims to provide the following research contributions:
  • We investigate how collaborative behavior in decentralized MAS varies across social choice methods, providing insights into how they influence overall cooperation and outcomes.
  • We identify key language features in multi-agent conversations as indicators of collaboration, offering a novel approach to analyzing linguistic cues in effective or ineffective interactions.
  • We propose various methods for determining the optimal stopping point in multi-agent collaboration, utilizing the linguistic features we identified.
2 Related Works
Multi Agent Frameworks
Recent advancements in MAS have led to the development of numerous platforms that facilitate collaboration among multiple agents. AutoGen (Wu et al., 2023) introduces a framework for building LLM-based applications where agents communicate with each other to accomplish tasks. CAMEL (Li et al., 2023a) is designed to enable autonomous collaboration among chat-based language models using role-playing and inception prompting. ChatDev (Qian et al., 2023) offers a chat-driven framework where LLM-powered agents work together to streamline software design, coding, and testing through unified language communication. MetaGPT (Hong et al., 2023) provides a meta-programming approach that encodes Standardized Operating Procedures into prompt sequences, facilitating efficient and error-reducing collaboration among LLM-based agents. AgentVerse (Chen et al., 2023) orchestrates collaborative groups of expert agents, drawing inspiration from human group dynamics to enhance task performance beyond the capabilities of individual agents. These MAS platforms address various aspects of the collaborative process, including agent profiling, recruitment, memory, planning, and communication protocols. However, they do not specifically investigate the group decision-making mechanisms within multi-agent interactions as we do in this paper. Some platforms employ centralized methods, and others only rely on unanimous or majority voting.
Multi-Agent Collaboration Environments
Multi-agent collaboration has been used to handle various applications. In robotics, there has been a continuous study regarding multi-agent collaboration in embodied environments (Dudek et al., 1996; Ota, 2006; Cena et al., 2013; Vorotnikov et al., 2018). In these studies, multi-agent collaboration in robotics has been explored through various frameworks, such as task allocation, coordination strategies, and communication protocols, to optimize collective performance and adaptability in dynamic environments. For LLM-based MAS, software development is a popular environment which inherently requires diverse expertise, continuous integration, and coordinated teamwork to build complex, functional systems efficiently (Qian et al., 2023; Hong et al., 2023; Dong et al., 2023). Additionally, LLM’s strong context understanding and immense parametric knowledge supported various world simulations with multi-agents. Including simulating society (Park et al., 2023; 2022), economy (Li et al., 2023b; Zhao et al., 2023), gaming (Xu et al., 2023; Wang et al., 2023), and psychology (Aher et al., 2023; Zhang et al., 2023b). These works require agents to interact with others in a group, and investigate their cooperate and competitive behavior.
3 Multi-Agent Collaboration and Group Decision-Making
In this section, we first introduce the formal definition of the setting this study uses for MAS, and propose RoundTable, a turn-based multi-agent collaboration platform for evaluating different decentralized group decision-making.
In a collaboration environment, there is a set of individuals (agents) $i\in{\mathbb{I}}$, and a set of all possible states of the world ${\mathbb{X}}=\{x_{1},x_{2},x_{3},...\}$. Then there is an individual utility function $u_{i}:{\mathbb{X}}\rightarrow{\mathbb{R}}$ for each agent that maps a state to a real number, representing preferences of agents. Based on the individual utility function, each agent proposes a state with proposal function $p_{i}=f_{p}(u_{i},C_{i},{\mathbb{X}}),p_{i}\in{\mathbb{X}}$, where $C_{i}$ is the context, including environment information, background, and collaboration history. Similarly, agent can also vote on proposals with voting function $v_{ij}=f_{v}(u_{i},C_{i},{\mathbb{S}}|p_{i}\in{\mathbb{S}}\subseteq{\mathbb{X}}),i,j\in{\mathbb{I}}$, where ${\mathbb{S}}$ represents a candidate list of proposals to vote on. With all proposals and agents’ votes, a social choice method (function) $F$ decides one group decision $x^{*}=F(p_{i},v_{ij}|i,j\in{\mathbb{I}})$. For example, when $F=Majority,p_{1}=Apple,p_{2}=Banana,p_{3}=Carrot,v_{1,j},v_{2,j}=(Yes,No,No),v_{3,j}=(No,Yes,No)$, then $x^{*}=p_{1}=Apple$.
3.1 RoundTable: Multi-Agent Collaboration Platform
With the definition above, we propose RoundTable, a multi-agent collaboration platform that can take various group decision-making mechanisms. The overview of RoundTable is shown in Figure 1. Due to the limited space, the details of the platform design and the LLM prompt are shown in Appendix A.1 and A.2. Following is a brief introduction of each phase.
  • Input and Initialization: RoundTable takes initial query (task), and initializes each agent giving agent specific background and utility function (or goals).
  • Collaboration Round: RoundTable employs round-based agent collaboration, with each round comprising three phases. Each phase occurs in a simultaneous open conversation, where agents act without order, and information is shared with everyone at the end of each phase. The process ends after $R$ rounds.
    • Message Phase: Each agent sends a message to its intended recipients.
    • Proposal Phase: Each agent proposes a potential solution, with an option to skip.
    • Voting Phase: Each agent votes for the candidate list, with an option to skip. The list consists with each agent’s latest proposal and latest accepted proposal by the group. With all votes, social choice method chooses the winner. If ties or disqualifies, the decision is deferred.
  • Output: After the final round, the latest accepted proposal will be selected as the final output of this system.
3.2 Social Choice Methods
To investigate the impact of different social choice methods on group decision-making in multi-agent collaboration, we compare the following 6 mechanisms:
@@ 3.2 Social Choice Methods
  • Rated Voting(Baujard et al., 2018): Each agent assigns ratings on a 5-point Likert scale to all candidate proposals, with 1 being the lowest and 5 being the highest. The proposal with the highest total score will be selected.
  • Ranked Voting(Arrow, 2012): Each agent ranks all candidate proposals from the most preferred to the least preferred. Social Choice will assign 1, 1/2, 1/3… points to the 1st, 2nd, 3rd… candidates on each ballot. The proposal with the highest total points will be selected.
  • Cumulative Voting(Black et al., 1958): For $|{\mathbb{I}}|$ candidate proposals, each agent is given $|{\mathbb{I}}|$ points to distribute among the proposals as they see fit. The proposal with the highest total points will be selected.
Unanimous, majority and plurality voting are one-vote mechanisms, where agents only choose the best candidate from the list; rated, ranked and cumulative voting are score-based mechanisms, which ask a nuanced, gradient preference over candidates.
@@ 4 Experiments
Using RoundTable, we explore multi-agent behavior patterns in $R=10$ rounds of collaboration. We evaluate RoundTable in two environments: simulated and complex.
4.1 Simulated Environment - Exchange Economy
@@ Introduction
We use an exchange economy for the simulated environment because its advantages make it well-suited for evaluating a decentralized MAS(Varian, 1992). First, it is a plus-sum game where agents seek equilibrium, meaning collaboration opportunities exist if allocation is not yet optimal. Second, multiple equilibria in the market allow dynamic collaboration. Third, while equilibrium doesn’t guarantee maximum utility, it must lie within one of the equilibria, helping to assess whether conflicts between agents hinder the group’s progress toward the ultimate goal.
Here, $K$ goods and $K$ agents participate in a market, with each good having a quantity of 100. Each agent has a Cobb-Douglas utility function $u_{i}=\prod a_{k}^{\theta_{k}},\sum\theta_{k}=1$, where $a_{k}$ represents the amount of good $k$ for agent $i$, and $\theta_{k}$ reflects the relative preference for each good (Cobb & Douglas, 1928). Agents aim to maximize their individual utility by finding an optimal allocation of goods. From a group perspective, the hidden goal is to maximize the total utility $U=\sum u_{i}$, which is not revealed to agents. For more on the exchange economy environment, see Appendix C.2.
To reflect common MAS collaboration patterns, we use an asymmetric utility function setup with $K=3$ agents, where each agent prefers different goods, mimicking real-world scenarios where agents specialize in different areas. Each agent’s utility function is $u_{i}=a_{i}^{0.8}\prod_{k\neq i}a_{k}^{\tilde{\theta}}$, where $\tilde{\theta}=\frac{1-0.8}{|{\mathbb{I}}|}$. Other types of utility function sets are shared in C.3. We report performance as the average of 100 simulations.
@@
We use various metrics to analyze multi-agent collaboration. For quality, we report the group total utility $U=\frac{\sum u_{i}}{U_{max}}$, and $U_{max}$ is the largest possible utility achievable in the environment. Efficiency is measured by the area under the curve (AUC), $AUC@n=\sum_{r=1}^{n}\frac{U_{r}}{U_{max}}$, where $r$ is round. Fairness is assessed by the Min/Max ratio, which compares the smallest individual utility to the largest at round 10: $\frac{u_{min}}{u_{max}}$. Rationality is the ratio of rounds where the proposed allocation’s utility exceeds the current one. [^1] This differs from compromise, where proposing lower utility than the previous proposal is natural, but a lower utility than the current allocation is irrational. Finally, rigidity measures how often an old allocation is kept, $rigidity=\frac{1}{10}\sum_{r=1}^{10}\bm{1}_{\mathrm{U_{r}=U_{r-1}}}$.
4.2 Complex Environment - Recommendation System
@@
For the complex environment, we focus on a recommendation system. It reveals unique collaborative behaviors absent in exchange economies. First, strong information asymmetry exists, as no single piece of information can accurately predict ratings. Second, group decision-making fosters diverse problem-solving approaches, with agents both consuming and contributing information in a feedback loop that enhances recommendations.
The task is to predict user ratings for specific items by analyzing historical interactions and identifying data patterns. This task is challenging for a single-agent system due to the overwhelming amount of scattered and limited essential information. In our setting, we divide the background information into several parts and assign each to different agents. The agents must then collaborate by exploring, analyzing, exchanging, and synthesizing their information to reach a comprehensive conclusion.
@@
We evaluate MAS performance using the MovieLens-100k dataset (Harper & Konstan, 2015). The dataset consists of three main tables: user, which contains demographic information of all users; item, which includes details of all movies, such as title, release date, and genre; and data, which holds 100,000 ratings from 943 users on 1,682 movies. Due to resource constraints, we randomly selected 100 examples from the u1.test split for evaluation. To simplify the data structure, we pre-processed it into three tables: basicInfo, userHistory, and movieHistory. Each table was assigned to a different agent, creating a 3-agent collaboration system. The definition and example of each table can be found in Appendix C.4.
@@
We use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to represent utility $U_{r}$, assessing the accuracy of predictions compared to the gold rating at round $r$. To evaluate MAS performance in the recommendation task, we compare it against two baselines: a simple Always Guess 4, which predicts the median regardless of input, and a strong State of the Art (SoTA) model from Behera & Nain (2023), which employs collaborative filtering with temporal features.
4.3 Cross-Agent Conversation Analysis
  • Information Difference measures how much new information is introduced. Effective collaboration should consistently bring new insights, while low information gain signals a stalled conversation. It is calculated as the average cosine distance between messages in the current round and the center embedding of the previous round.
  • Dialogue Acts are communicative functions that capture actions, intents, and behaviors within messages as discrete states. We design a set of dialogue acts for multi-agent collaboration: Inform, Request, Confirm, Summarize, Evaluate, Propose, Compromise, Defend, Accept, Decline, and use LLM-labeling to perform multi-label classification on a target message based on the previous round. Details of the dialogue act labels and prompt can be found in Appendix C.5.
4.4 Dialogue Act Transition Graph
We constructed transition graphs to visualize the dynamics of dialogue act transitions across rounds, highlighting the most frequent transitions. Each node represents a dialogue act, and directed edges indicate the probability of transitions between acts from one round to the next. The figure displays only the most probable edge per node, excluding self-loops. The detailed definition of nodes and edges is in Appendix C.6.
4.5 Early Stopping in MAS
In the previous experiments, we enforced 10 rounds of collaboration to compare agent behaviors and identify patterns. However, effective early stopping can prevent redundancy, avoid stagnation, and optimize decision-making by reducing unnecessary iterations. In this experiment, we evaluate the following early stopping methods:
  • First Agreement stops collaboration whenever a proposal passes the social choice criterion.
  • Information Difference is the average embedding distance captured at the round that produced the best outcomes in the train set, and stops collaborations in the test set when the distance became lower than the threshold. This idea is supported by the observation in Section 5.3.1.
  • Dialogue Act method utilizes pairs of dialogue acts, linking one from the previous round to another from the current round. We perform ordinary least squares (OLS) regression on all such pairs in relation to the performance. The regression coefficients indicate the most impactful pairs that act as stop signals. Further details of the algorithm are provided in Appendix C.7.

### v2

@@ Abstract
Effective group decision-making is critical in Multi-Agent Systems (MAS). Yet, how different mechanisms for reaching consensus impact collaboration quality and efficiency remains understudied. We conduct a systematic study on group decision-making mechanisms in a decentralized setting. Through controlled experiments, we analyze how different voting rules affect decision quality and efficiency in a multi-round collaboration. Results reveal that majority voting often cause inefficient collaboration due to its strict acceptance criteria. At the extreme, unanimous voting gives 87% lower initial performance than the best-performing method. Our qualitative analysis of cross-agent communication shows that messages become longer and more repetitive over time: while message length increases by 84%, similarity to the previous round increases to 90%. Based on these insights, language-based early stopping methods make the performance 13% closer to oracle while reducing rounds by 50%. Our findings highlight the crucial role of group decision-making in optimizing MAS collaboration.
RoundTable: Investigating Group Decision-Making Mechanism in Multi-Agent Collaboration
Young-Min Cho1,2thanks: Work completed during employment at Amazon., Raphael Shu2, Nilaksh Das2, Tamer Alkhouli2 Yi-An Lai2, Jason Cai2, Monica Sunkara2, Yi Zhang2, Dan Roth1* 1University of Pennsylvania, 2AWS GenAI jch0@seas.upenn.edu {zhongzhu, yizhngn}@amazon.com
@@ 1 Introduction
Collaboration is a fundamental aspect of human society. It enables individuals to address complex challenges that would be unmanageable otherwise (De Man and Duysters, 2005; Graesser et al., 2018; Bittner and Leimeister, 2013). Inspired by this principle, Large Language Model (LLM)-based Multi-Agent Systems (MAS) have emerged as a paradigm where autonomous LLM agents interact, reason, and coordinate to contribute diverse perspectives, fostering richer problem-solving through collective intelligence. This approach is valuable in complex domains such as software development Hong et al. (2023); financial trading, Li et al. (2023); and healthcare Tang et al. (2023). These systems demonstrate the importance of MAS, which coordinates interactions among agents to enhance task performance and solution quality.
Figure 1: An example of decentralized multi-agent collaboration, in which agents reach consensus through multi-round communication and decision-making.
One fundamental component in MAS is decision-making mechanism, which defines how decisions are made in the system. It enables agents to coordinate their actions and achieve collective goals efficiently. Decision-making can be categorized into two levels: 1) individual decision-making, where each agent independently evaluates its environment, sets objectives, and selects actions to optimize its own outcomes; and 2) group (collective) decision-making, where a framework aggregates the opinions of all agents to produce a collective outcome, with implicit early stopping when consensus stabilizes. Since agents likely to have varying perspectives by the nature of MAS, aggregating different opinions to make a group level decision becomes especially important.
Although individual decision-making is mostly powered by LLM, group decision-making varies across systems. While centralized MAS resolving conflicts between agents through a leader or hierarchical decision-making pipeline (Hong et al., 2023; Qian et al., 2023; Dong et al., 2023), decentralized MAS distributes power equally among agents (Mandi et al., 2024; Zhang et al., 2023; Xu et al., 2023; Park et al., 2023). Despite the recent success of decentralized MAS, such as Multi-Agent Debate Du et al. (2023); Chan et al. (2023); Liang et al. (2023), existing studies predominantly rely on majority voting or average rating without thoroughly evaluating alternative mechanisms.
Indeed, the aggregation of diverse opinions has been extensively studied in economics and political science. As a theoretical framework for decision-making, social choice theory examines how collective decisions are made by aggregating individual preferences into a unified outcome Arrow (2012); Riker (1988); Endriss (2008). Within social choice theory, voting rules are the most representative and widely used functions that maps individual votes to a collective outcome. Examples of voting rules include unanimous, majority, plurality, rated, ranked, and cumulative voting.
While traditional studies in economics and computational social choice are extensive, investigating social choice in LLM-based multi-agent systems remains non-trivial. This is because the mere acknowledgment of using a specific voting rule can influence agents’ actions and strategies. In decentralized MAS, where agents are adaptive and capable of strategic behavior, the voting rule is not just an aggregation tool but also a factor that shapes the dynamics of interaction. As shown in Figure 1, this feedback loop between the voting rule and agent behavior makes it essential to investigate how different voting rules affect not only the final outcomes but also the process of reaching consensus. Furthermore, the decision-making mechanism requires a stopping signal during collaboration to reduce communication costs and enhance collaboration quality. However, social choice theory does not account for communication dynamics or provide criteria for early stopping. Therefore, studying the decision-making mechanisms in MAS is critical for designing systems that ensure robust, fair, and efficient collective intelligence.
This study investigates group decision-making mechanisms in decentralized MAS. We develop an experimental platform, RoundTable, which enables multi-round, multi-agent decentralized collaboration and allows controlled experiments to compare the effects of different decision-making mechanisms. By comparing various voting rules, we see that too strict voting threshold and format yield worse performance. Unanimous voting, which has the strictest threshold and format, shows 87% lower initial round performance and 39% lower final round performance. We further analyze cross-agent communications to characterize MAS interactions, finding that messages grow longer and more repetitive over time. Message length increases by 84%, while similarity to the previous round reaches 90%, indicating inefficiencies in prolonged collaboration. Based on these insights, we find language-based early stopping improves performance by 13% while halving the number of rounds, enabling timely, efficient collaboration.
This paper provides the following research contributions:
  • We show that majority voting often fails in multi-agent systems because its strict acceptance criteria hinder decision-making.
  • We identify that extended collaboration often leads to inefficiencies, characterized by longer and more repetitive messages, which result in redundant communication and reduced effectiveness.
  • Our experiments show that can be leveraged for early stopping, enhancing performance and reducing the number of rounds.
2 Related Work
Centralized vs Decentralized MAS
Centralized MAS have been extensively studied, with systems like MetaGPT (Hong et al., 2023) and AutoGen (Wu et al., 2023) using manager agents or hierarchical structures for decision-making. While effective for coordination, these systems face issues such as single points of failure and decision biases due to concentrated authority (Owens et al., 2024; Jiang and Lu, 2019).
Decentralized MAS, as seen in RoCo (Mandi et al., 2024) and Generative Agents (Park et al., 2023), distribute decision power among agents, enhancing flexibility, robustness, and adaptability. Despite these strengths, current research often overlooks challenges like consensus-building and efficient cross-agent communication, especially in complex, dynamic environments (Xu et al., 2023; Zhang et al., 2023).
Social Choice Theory
Social choice theory, rooted in economics, examines how individual preferences can be aggregated for collective decision-making, with foundational work like Arrow’s Impossibility Theorem highlighting the challenges of fairness (Arrow, 2012). It has since expanded into computational social choice for algorithmic decision-making and automated negotiation to resolve conflicts in multi-agent settings (Endriss, 2008; Kraus, 2001). In LLM-based MAS, voting rules help consolidate diverse agent opinions, but most applications rely heavily on majority voting due to its simplicity (Li et al., 2022; Chan et al., 2023). This over-reliance limits adaptability, overlooking more flexible voting rules like ranked or rated voting that could improve collaboration and decision efficiency in dynamic MAS environments.
Cross-Agent Communication
Cross-agent communication is essential in decentralized MAS, enabling information sharing, negotiation, and consensus-building without central control (Jennings et al., 2001; Park et al., 2023; Han et al., 2024). However, structured protocols like turn-taking and fixed roles can create power imbalances, undermining decentralization. While prior work optimizes message passing and protocols, how to use language patterns to understand collaboration remains underexplored. This study examines language cues as indicators of collaboration quality and introduces language-based early stopping to enhance decision-making efficiency.
3 Decentralized MAS
A decentralized MAS is a collection of autonomous agents collaborating towards a common goal without a central controller. Each agent in the system operates based on its individual context, defined by available information, historical interactions, and personalized objectives. The collaboration unfolds over multiple rounds, where agents communicate, propose solutions, and vote to reach a collective decision. This system is characterized by 1) The collective decision emerges from individual proposals and votes rather than being imposed by a single entity; 2) Agents update their strategies based on past interactions, ensuring adaptability to changing circumstances.
@@ 3 Decentralized MAS
Figure 2: Overview of our multi-agent collaboration platform: RoundTable. It allows a multi-round collaboration where agents simultaneously send messages, propose solutions, and vote. Based on a voting rule, RoundTable selects the most preferred proposal for group decisions. Detailed introduction of RoundTable is in Appendix B.1.
3.1 Problem Definition
We consider an environment where all possible states of the world are defined as $\mathcal{X}$ and a task $\Sigma$ is set as the global objective. Here, a decentralized MAS system is defined as a collaborative decision-making process over $R$ rounds, involving a set of agents $i\in\mathcal{I}$.
Agent Definition
At each round $r$, an agent $i$ takes actions with LLM and its latest individual context, denoted as $\Omega_{i}=\{B_{i},H\}$.
• Background $B_{i}$ includes: Task $\Sigma$, Collaboration Rules, and Self Information (persona, data access, and objective).
• Latest History $H$ includes: Messages $M$, Proposals $P$, Votes $V$, and Group Decision $x^{*}_{r}\in\mathcal{X}$: The latest decision reached by the group at round $r$.
Action Definition
The collaboration at each round $r$ is structured into three ordered phases, which defines action space for agents:
• Message Phase: Each agent formulates a message based on its context: $m_{i,r}=f_{m}(\Omega_{i})$
• Proposal Phase: Each agent generates a proposal for a potential group decision: $p_{i,r}=f_{p}(\mathcal{X},\Omega_{i}),\quad p_{i,r}\in\mathcal{X}$
• Voting Phase: Each agent evaluates and votes on proposals: $v_{i,p,r}=f_{v}(p,\Omega_{i})$, where the format of vote is defined by the input voting rule $F$.
Voting Rules
At the end of each round, a voting rule $F$ determines the collective decision for the round:
$\hat{x^{*}}_{r}=F(\{(c,\{v_{i,c,r}\mid i\in\mathcal{I}\})\mid c\in C\})$
where the candidate set $C$ consists of all proposals and the previous group decision: $C=P\cup\{\hat{x^{*}}_{r-1}\}$.
To investigate the impact of different voting rules in multi-agent collaboration, we compare the following six voting rules: Unanimous, majority and plurality voting, where agents only choose the best candidate from the list (discrete voting rules); rated, ranked and cumulative voting, which ask a nuanced, gradient preference over candidates (preference voting rules). [^1] Mathematical definitions and prompt texts are presented in Appendix C.
@@ 3.1 Problem Definition
• Rated Voting (Baujard et al., 2018): Each agent rates proposals on a 5-point scale, selecting the one with the highest total score.
• Ranked Voting (Arrow, 2012): Agents rank proposals by preference, with points assigned as 1, 1/2, 1/3, etc., and the highest-scoring proposal is selected.
• Cumulative Voting (Black, 1958): Each agent distributes $|\mathcal{I}|$ points among $|\mathcal{I}|$ proposals, selecting the one with the highest total.
The goal of the decentralized MAS is to find the solution $x^{*}\in\mathcal{X}$ with inputs: the task $\Sigma$, all agent information $\{\phi_{i}|\forall i\in\mathcal{I}\}$, and a voting rule $F$. Based on the definitions above, we propose RoundTable, an experiment platform for decentralized multi-agent collaboration that can compare various voting rules through controlled experiments. The overview of RoundTable is shown in Figure 2.222details of the platform design and the LLM prompts are shown in Appendix B.1 and B.2.
@@ 4 Experiments
Using RoundTable, we analyze multi-agent behavior patterns over $R=10$ rounds per collaboration. Since disagreement among agents creates tension between cooperation and competition, we evaluate RoundTable in two environments: competitive-focused and cooperative-focused. In both, we report performance as the average of 100 independent collaborations.
4.1 Competitive Environment - Exchange Economy
@@ Introduction
We use an exchange economy for the competitive environment as the advantages make it ideal for evaluating a decentralized MAS (Varian, 1992). First, it is a plus-sum game where agents seek equilibrium, meaning collaboration opportunities exist if allocation is not yet optimal. Second, multiple equilibria in the market allow dynamic collaboration. Third, while equilibrium doesn’t guarantee maximum utility, it must lie within one of the equilibria, helping to assess whether conflicts between agents hinder the group’s progress toward the ultimate goal.
Here, $K$ types of goods and the same number of agents participate in a market, with each good having a quantity of 100. Each agent has a Cobb-Douglas utility function $u_{i}=\prod a_{k}^{\theta_{k}},\sum\theta_{k}=1$, where $a_{k}$ represents the amount of good $k$ for agent $i$, and $\theta_{k}$ reflects the relative preference for each good (Cobb and Douglas, 1928). Agents’ objective is set to maximize their individual utility by finding an optimal allocation of goods. From a group perspective, the hidden goal is to maximize the total utility $U=\sum u_{i}$, which is not revealed to agents. [^3] For more on the exchange economy environment, see Appendix D.1.
To reflect common MAS collaboration patterns, we use an asymmetric utility function setup with $K=3$ types of goods and agents, where each agent prefers different goods, mimicking real-world scenarios where agents specialize in different areas. [^4] Other types of utility function sets are shared in Appendix D.2. Each agent’s utility function is $u_{i}=a_{i}^{0.8}\prod_{k\neq i}a_{k}^{\tilde{\theta}}$, where $\tilde{\theta}=\frac{0.2}{K-1}$.
@@
We use various metrics to analyze multi-agent collaboration. For quality, we report the group total utility $U=\frac{\sum u_{i}}{U_{max}}$, where $U_{max}$ is the largest possible utility achievable in the environment. Efficiency is measured by the area under the curve (AUC), $AUC@n=\sum_{r=1}^{n}\frac{U_{r}}{U_{max}}$. Fairness is assessed by $\frac{u_{min}}{u_{max}}$, which compares the smallest individual utility to the largest at the last round. Rationality is the ratio of rounds where the individual utility of the proposed allocation exceeds the current one. [^5] This differs from compromise, where proposing lower utility than the previous proposal is natural, but a lower utility than the current allocation is irrational. Finally, rigidity measures how often an old allocation is kept, $rigidity=\frac{1}{10}\sum_{r=1}^{10}\mathbbm{1}(U_{r}=U_{r-1})$.
4.2 Cooperative Environment - Recommendation System
@@
For the cooperative environment, we focus on the recommendation system. It reveals unique collaborative behaviors absent in exchange economies. First, strong information asymmetry exists, as no partial information can accurately predict. Second, group decision-making fosters diverse problem-solving approaches, with agents both consuming and contributing information in a feedback loop that enhances recommendations.
The task is to predict user ratings for specific items by analyzing historical interactions and detecting trends and correlations in the data. This is challenging for a single-agent system due to the overwhelming amount of data, but essential information is scattered and limited. In our setting, we divide the background information into three parts and assign each to different agents. The agents must then collaborate by exploring, analyzing, exchanging, and synthesizing their information to reach a comprehensive conclusion.
@@
We evaluate MAS performance using the MovieLens-100k dataset (Harper and Konstan, 2015). The dataset consists of information of all users and movies, and the rating history of users on movies. Due to resource constraints, we randomly selected 100 examples from the u1.test split for evaluation. To simplify the data structure, we pre-processed it into three tables: basicInfo, userHistory, and movieHistory. Each table was assigned to a different agent, creating a 3-agent collaboration system. [^6] The definition and example of each table can be found in Appendix D.3.
@@
We use Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) to represent utility $U_{r}$, assessing the accuracy of predictions compared to the gold rating at round $r$. To evaluate MAS performance in the recommendation task, we compare it against two baselines: a simple Always Guess 4, which predicts the median regardless of input, and a strong State of the Art (SoTA) model from Behera and Nain (2023), which employs collaborative filtering with temporal features.